OpenStudy (timinthor):

PLEASE HELP! WILL MEDAL! Transform this equation to slope-intercept form: 6x + 4y = 8

OpenStudy (texaschic101):

slope intercept form is y = mx + b 6x + 4y = 8 ---subtract 6x from both sides 4y = -6x + 8 -- now divide both sides by 4 y = -6/4x + 8/4 -- reduce y = - 3/2x + 2

OpenStudy (timinthor):

Thank you @texaschic101